# Diversity Methods for Improving Convergence and Accuracy of Quantum Error Correction Decoders Through Hardware Emulation

### v1

@@ 6 Conclusion
We have designed a hardware emulator implemented on an FPGA that enables us to explore, within days, the lower logical error rate region necessary for implementing fault-tolerant quantum computation. This architecture is sufficiently flexible to run real QEC architectures while also monitoring and storing various parameters and error patterns. The platform is vendor-independent, and the results obtained have been instrumental in designing diversity-based decoding methods to enhance both convergence and accuracy. These diversity methods are based on analyzing and leveraging quantization noise, as well as employing various low-latency BP implementations.
We analyze the results of both diversity methods under phenomenological and circuit-level noise, respectively, demonstrating significant improvements in speed and accuracy. These examples illustrate the value of examining the decoding problem from a hardware perspective, through hardware emulation, considering the specific characteristics of the architectures as part of the co-design process.
Future work will focus on analyzing both diversity proposals in tandem, combining the diversity based on quantization noise with BP implementations to further reduce the number of LSD or OSD executions. Additionally, we will investigate the behavior of LSD and OSD after filtering a larger number of error patterns through diversity, with the goal of simplifying the algorithms to create more efficient and scalable implementations.
@@ 7 Acknowledgment
This work was supported by the QuantERA grant EQUIP (Spain MCIN/AEI/10.13039/501100011033, grant PCI2022-132922), funded by Agencia Estatal de Investigación, Ministerio de Ciencia e Innovación, Gobierno de España and by the European Union “NextGenerationEU/PRTR". This research is part of the project PID2023-147059OB-I00 funded by MCIU/ AEI/ 10.13039/501100011033/ FEDER, UE. F. Garcia-Herrero’s work on this project was partially funded by a grant from Google Quantum AI.

### v2

@@ 5 Conclusion
We have designed a hardware emulator implemented on an FPGA that enables us to explore, within days, the lower LER region necessary for implementing fault-tolerant quantum computation. This architecture is sufficiently flexible to run real QEC architectures while also monitoring and storing various parameters and error patterns. The platform is vendor-independent, and the results obtained have been instrumental in designing diversity-based decoding methods to enhance both convergence and accuracy. These diversity methods are based on analyzing and leveraging quantization noise, as well as employing various low-latency BP implementations.
We analyze the results of both diversity methods under code-capacity and circuit-level noise, respectively, demonstrating significant improvements in speed and accuracy. These examples illustrate the value of examining the decoding problem from a hardware perspective, through hardware emulation, considering the specific characteristics of the architectures as part of the co-design process.
Future work will focus on analyzing both diversity proposals in tandem, combining the diversity based on quantization noise with BP implementations to further reduce the number of LSD or OSD executions. Additionally, we will investigate how the diversity approach can be integrated with other ensemble-based methods, including Relay-BP, to reduce the number of BP decoders required while maintaining or improving performance. We also plan to explore the behavior of LSD and OSD after filtering a larger number of error patterns through diversity, with the goal of simplifying these algorithms and making them more efficient and scalable for future quantum computing applications.
@@ 6 Acknowledgment
This work was supported by the QuantERA grant EQUIP (Spain MCIN/AEI/10.13039/501100011033, grant PCI2022-132922), funded by Agencia Estatal de Investigación, Ministerio de Ciencia e Innovación, Gobierno de España and by the European Union “NextGenerationEU/PRTR”. This research is part of the project PID2023-147059OB-I00 funded by MCIU/ AEI/ 10.13039/501100011033/ FEDER, UE. F. Garcia-Herrero’s work on this project was partially funded by a grant from Google Quantum AI.